OpenStudy (anonymous):

A basketball team has 8 players. In how many different ways can the coach select 2 players to be the captains for tonight's game? A. 16 B. 28 C. 56 D. 64

OpenStudy (lieutenantgeneral):

8^2 = 64 8 * 2 = 16

OpenStudy (anonymous):

28 is the answer

OpenStudy (igreen):

\(\dfrac{8 \cdot 7}{2}\) \(\dfrac{56}{2}\) \(28\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @leena_paige2

OpenStudy (igreen):

@danish071996 You shouldn't give out direct answers without explanation.

OpenStudy (anonymous):

okk bt i knew u will explain her